# Exploring Persona Sentiment Sensitivity in Personalized Dialogue Generation

### v1

@@ Acknowledgement
This research was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).
Appendix A Which Factors Influence the Polarity-Based Trend?
@@
A.1 Varying Persona Count
@@
A.2 Impact of Polarity Ratio
@@ A Notable Anomaly in Mixed Ratios
A.3 Effect of Model Size
@@
To determine whether polarity-based trends remain consistent across different model sizes, we employ various Qwen-2.5 models ranging from 0.5B to 3B, 7B, 14B, and 32B, alongside the default 7B model. For practical purposes, we sample 1K user pairs from the pool using the Negative Pairing, Positive Pairing, and Mixed Pairing strategies in §3, ensuring identical inputs for each model. The dialogue metrics follow those in AppendixA.1.
@@
Appendix B Evaluation
B.1 Consistency
@@ B.2 Coherence
Appendix C Prompt Templates
@@
Appendix D Dialogue Statistics
@@
Appendix A Negative pairings (K=1) 500 13.61 10.29 281.76

### v2

@@ Acknowledgement
This research was supported by the Chung-Ang University Graduate Research Scholarship in 2025. This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)].
Appendix A Why Are Large Language Models Sensitive to Persona Sentiment?
We have observed the performance gap across different persona polarities in §3. As we utilize an LLM without any additional fine-tuning, we could infer that LLMs inherently exhibit sensitivity to persona sentiment. Consequently, we explain that this sensitivity is mainly due to the scarcity of negative expressions in the pretraining data, which may have been intentionally limited to prevent toxicity. And Figure 2 clearly illustrates a lower proportion of negative personas compared to other polarities. Such data imbalance is evident even in widely-used foundational datasets like ConvAI2, suggesting similar limitations in collecting sufficient negative persona data in other datasets or real-world contexts. Additionally, we also find that LLM’s sensitivity can stem from post-training techniques that steer models toward generating responses preferred by humans. For the experiment, we sample 1,000 dialogues per model under both the negative and positive configurations described in §3.1.1. For the neutral configuration, we follow a similar approach by grouping five intermediate-stage personas from Figure 3 into pairs, generating 1,000 dialogues for each model accordingly.
Pairing Persona Dialogue
Negative 1.71 2.67 (+0.96)
Positive 4.25 4.52 (+0.27)
Neutral 3.06 4.09 (+1.03)
Table 5: Average sentiment scores (1–5 Likert) for dialogues and their personas, rated by GPT-4o. Lower scores denote negativity, higher positivity; red values mark the dialogue’s sentiment gain over its persona.
As shown in Table 5, regardless of model type, we observe a consistent tendency to produce more positive dialogues than the assigned personas. Models consistently generate more positive dialogues, with the gap widening when dialogue quality decreases. This trend suggests that post-training practices bias models toward positive outputs, leading to confusion when persona sentiment diverges significantly, lowering dialogue quality.
Appendix B Which Factors Influence the Polarity-Based Trend?
@@
B.1 Varying Persona Count
@@
B.2 Impact of Polarity Ratio
@@ A Notable Anomaly in Mixed Ratios
B.3 Effect of Model Size
@@
To determine whether polarity-based trends remain consistent across different model sizes, we employ various Qwen-2.5 models ranging from 0.5B to 3B, 7B, 14B, and 32B, alongside the default 7B model. For practical purposes, we sample 1K user pairs from the pool using the Negative Pairing, Positive Pairing, and Mixed Pairing strategies in §3, ensuring identical inputs for each model. The dialogue metrics follow those in AppendixB.1.
@@
Appendix C GPT-4o Results
We extended our analysis by incorporating GPT-4o—the leading closed-source LLM in many chatbot benchmarks—into our backbone set to test whether the phenomenon identified in RQ1 persists. The experimental setup mirrors that described in §3.1.1.
Pairing Consistency Coherence
C score $\uparrow$ Contd. $\downarrow$ G-eval $\uparrow$ PairEval $\uparrow$ G-eval $\uparrow$
Original 0.385 7.92 4.14 2.79 4.56
Negative 0.626 11.11 4.27 2.75 4.44
Positive 0.651 4.12 4.42 2.83 4.69
Mixed 0.638 8.60 4.32 2.76 4.54
Opposite 0.589 8.64 4.23 2.75 4.44
Table 6: Combined consistency and coherence results of the dialogues generated by GPT-4o.
As shown in Table 6, the same pattern emerges with GPT-4o: the positive-pairing configuration yields the highest dialogue quality across all metrics, whereas configurations containing negative pairings or mixed sentiment polarities perform noticeably worse. In particular, the original-pairing configuration—which reflects real-world distributions—suffers a larger quality drop than we observed with open-source LLMs. This result reinforces the robustness of our findings and warns of potential, unforeseen performance degradation in downstream applications.
Appendix D Evaluation
D.1 Consistency
@@ D.2 Coherence
Appendix E Prompt Templates
@@
Appendix F Dialogue Statistics
@@
Appendix B Negative pairings (K=1) 500 13.61 10.29 281.76